OpenStudy (anonymous):

find the x-intercepts and y intercepts f(x)=-x^2+2x+24

OpenStudy (turingtest):

x-intercepts\[0=-x^2+2x+24=(4+x)(6-x) \to x=\left\{ -4,6 \right\}\]y-intercepts \[y=-(0)^2+2(0)+24=24\]